OpenStudy (anonymous):

Given a Poisson process with an average, a arrivals per unitof time, find the probability that there will be no arrivals during a time interval of length, t, namely, the probability that the waiting times between successive arrivals will be at least of length t.

OpenStudy (anonymous):

I think I have to use the gamma distribution formula here

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ \beta^{\alpha}\Gamma(\alpha) }x^{\alpha-1}e^{-x/\beta}\] for x>0, alpha>0,Beta>0